# Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models

### v4

@@ B.2 Proof of Lemma 2
To complete the first part of the proof, it needs to be shown that the diagonal elements of $D$ are either $1$ or $-1$. By the normalization of the unit variances of $e_{t}$ and $\tilde{e}_{t}$, $\text{Cov}(\tilde{e}_{t}|\mathcal{F}_{t-1})=\text{Cov}(Ce_{t}|\mathcal{F}_{t-1})=I_{d}$, which is equivalent to $C\text{Cov}(e_{t}|\mathcal{F}_{t-1})C^{\prime}=I_{d}$ and thereby $CC^{\prime}=I_{d}$. By substituting $C=DP$ to this expression, we obtain $DPP^{\prime}D^{\prime}=I_{d}$ and by the orthogonality of permutation matrix $DD^{\prime}=I_{d}$. That is, $d_{ii}^{2}=1$ for all $i=1,...,d$, so $d_{ii}=\pm 1$. Since diagonal elements of $D$ are, hence, either $1$ or $-1$, it follows that $B_{y,t}$ is unique up to ordering of its columns and changing all signs in a column almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
Threshold weights. To show the ”moreover”-part, consider first the threshold transition weights defined in (2.3). Denote the partition of the time periods to the regimes as $\mathcal{T}_{m}=\{t:r_{m-1}<z_{t}\leq r_{m}\}$, $m=1,...,M$, where $z_{t}$ is the switching variable and $r_{0}=-\infty,r_{M}=\infty$, and $r_{1},...,r_{M-1}$ are threshold parameters, and suppose Equation (B.2) holds also for some other partition $\tilde{\mathcal{T}}_{m}=\{t:\tilde{r}_{m-1}<z_{t}\leq\tilde{r}_{m}\}$, $m=1,...,M$. By Condition (3) of Lemma 2, the sets $\mathcal{T}_{m}$ and $\tilde{\mathcal{T}}_{m}$ are nonempty for all $m=1,...,M$. That is, we assume Equation (B.4) holds for the two sets of parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,...,M$, $r_{1},...,r_{M-1}$ and $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,...,M$, $\tilde{r}_{1},...,\tilde{r}_{M-1}$.
At each $t$, the model reduces to a linear VAR $y_{t}=\phi_{m}+\sum_{i=1}^{p}A_{m,i}y_{t-1}+B_{m}e_{t}$ corresponding to the regime $m$ for which $r_{m-1}<z_{t}\leq r_{m}$. Consequently, Equation (B.4) can be written as
$\tilde{\phi}_{\tilde{m}_{t}}-\phi_{m_{t}}+\sum_{i=1}^{p}(\tilde{A}_{\tilde{m}_{t},i}-A_{m_{t},i})y_{t-i}=0,$ (B.6)
where $m_{t}\in\{1,...,M\}$ is the regime $m$ that satisfies $r_{m-1}<z_{t}\leq r_{m}$ and $\tilde{m}_{t}\in\{1,...,M\}$ is the regime $m$ that satisfies $\tilde{r}_{m-1}<z_{t}\leq\tilde{r}_{m}$ at the time $t$.
Equation (B.6) defines $d$ scalar equations such that the left side of each scalar equation is a first degree polynomial in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$, where, for the $j$th equation, the coefficients are given by the $j$th element of $\tilde{\phi}_{\tilde{m}_{t}}-\phi_{m_{t}}$ and the elements of the $j$th rows of $\tilde{A}_{\tilde{m}_{t},i}-A_{m_{t},i}$, $i=1,...,p$. Hence, by applying Lemma B.1 to each scalar equation in (B.6), it can be concluded that the set of solutions to Equation (B.6) has a Lebesgue measure zero in $\mathbb{R}^{dp}$ unless it is identically true. Since Equation (B.6) is identically true if and only if $\tilde{\phi}_{\tilde{m}_{t}}=\phi_{m_{t}}$ and $\tilde{A}_{\tilde{m}_{t},i}=A_{m_{t},i}$, $i=1,...,p$, and because Equation (B.6) must hold for all $t$ (thereby for all regimes), it follows that $\tilde{\phi}_{\tilde{m}}=\phi_{m}$ and $\tilde{A}_{\tilde{m},i}=A_{m,i}$, $i=1,...,p$, for all $m=1,...,M$ almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ (as a countable union of measure zero sets has a measure zero). It then follows from Condition (2) of Lemma 2 that once $\tilde{\phi}_{\tilde{m}}=\phi_{m}$ and $\tilde{A}_{\tilde{m},i}=A_{m,i}$, $i=1,...,p$, for all $m=1,...,M$, the only way for these parameters to match the data almost everywhere is that the same regime index must apply for each $t$. Consequently, $\tilde{m}_{t}=m_{t}$ for all $t$, so the threshold parameters $r_{1},...,r_{M-1}$ and $\tilde{r}_{1},...,\tilde{r}_{M-1}$ imply identical partitions of the data to the regimes, $\mathcal{\tilde{T}}_{m}=\mathcal{T}_{m}$ for all $m=1,...,M$ almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.
Because a regime switches when (and only when) the switching variable passes one of the threshold parameter values, for a given (countable number of) observations of $z_{t}$, the threshold parameters can only be identified with the resolution determined by the spacing of the order statistics of $z_{t}$. For instance, if the $n$th ($n\in\{1,...,T-1\}$) order statistic of $z_{t}$ is $1.00$, its $n+1$th order statistic is $1.02$, and $M=2$, then all values of the threshold parameter $r_{1}$ between $1.00$ and $1.02$ correspond to the same model, as they all imply the same regime-switches. Hence, the identification of the threshold parameters should be understood in the sense that they are identified to be between some order statistics of the switching variable series $z_{t}$. Now, if the threshold parameter values $\tilde{r}_{1},...,\tilde{r}_{M-1}$ are between different order statistics of $z_{t}$ than $r_{1},...,r_{M-1}$, the partitions $\mathcal{\tilde{T}}_{m}$ and $\mathcal{T}_{m}$ are different for some $m\in\{1,...,M\}$, in which case, by the above discussion, Equation (B.6) can only hold in sets of Lebesgue measure zero in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$. This concludes that the parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,...,M$ and $r_{1},...,r_{M-1}$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.
Logistic weights. Next, we show the ”moreover” part for the logistic transition weights defined in (2.4) with $M=2$. Suppose Equation (B.4) holds for the two sets of parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,...,M$, $c,\gamma$ and $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,...,M$, $\tilde{c},\tilde{\gamma}$. The proof proceeds in the following steps: (1) we show that $\tilde{\alpha}_{m,t}=\alpha_{m,t}$ for each $t$ almost everywhere in $z_{t}\in\mathbb{R}$, (2) we use this identity to show that the parameters $c,\gamma$ are uniquely identified (almost everywhere), and then (3) we use these results to show that the intercepts and autoregression matrices are uniquely identified.
By applying Lemma B.1 to each scalar equation in Equation (B.4), it follows that Equation (B.4) holds almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ if only if Equation (B.4) is identically true, that is, $\tilde{\phi}_{y,t}=\phi_{y,t}$ and $\tilde{\boldsymbol{A}}_{y,t}=\boldsymbol{A}_{y,t}$ (irrespective of the weight function dynamics). By making use of the identity $\alpha_{1,t}=1-\alpha_{2,t}$, these equations can be written as
$\tilde{\phi}_{1}-\phi_{1}+\alpha_{2,t}(\phi_{1}-\phi_{2})-\tilde{\alpha}_{2,t}(\tilde{\phi}_{1}-\tilde{\phi}_{2})=0$ (B.7)
and
$\tilde{A}_{1,i}-A_{1,i}+\alpha_{2,t}(A_{1,i}-A_{2,i})-\tilde{\alpha}_{2,t}(\tilde{A}_{1,i}-\tilde{A}_{2,i})=0,\ \ i=1,...,p,$ (B.8)
respectively.
Equations (B.7) and (B.8) define a system of $d(1+pd)$, $d\geq 2,p\geq 1$, linear scalar equations in two variables $\alpha_{2,t}$ and $\tilde{\alpha}_{2,t}$ (interpreting the AR parameters as coefficients here). Suppose first that all of them are identically true, i.e., that all the coefficients are zero. In (B.7), this requires that $\tilde{\phi}_{1}=\phi_{1}$, $\phi_{1}=\phi_{2}$, and $\tilde{\phi}_{1}=\tilde{\phi}_{2}$. Similarly in (B.8), zero coefficients require that $\tilde{A}_{1,i}=A_{1,i}$, $A_{1,i}=A_{2,i}$ and $\tilde{A}_{1,i}=\tilde{A}_{2,i}$ for all $i=1,...,p$. But as $\phi_{1}=\phi_{2}$ and $A_{1,i}=A_{2,i}$ for all $i=1,...,p$ violates Condition (2) of Lemma 2, it must be that not all scalar equations in (B.7) and (B.8) are identically true.
For each scalar equation that is not identically true, the coefficients are not all zero, and each of these scalar equations can be expressed in the form $a_{2,j}\alpha_{2,t}=a_{0,j}+a_{3,j}\tilde{\alpha}_{2,t}$ for some $a_{0,j},a_{2,j},a_{3,j}\in\mathbb{R}$ for the $j$th not identically true equation. If $a_{2,j}=0$, then $\tilde{\alpha}_{2,t}$ is the constant $-a_{0,j}/a_{3,j}$ for all $t$, which violates Condition (3) of Lemma 2. Hence, $a_{2,j}\neq 0$ (in the not identically true equations), implying that $\alpha_{2,t}=a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}$, where $a_{1,j}\equiv a_{3,j}a_{2,j}^{-1}$. Thus, the solution to Equations (B.7) and (B.8) must satisfy this relationship (for all $j$).
Making the dependence of the transition weights of $z_{t}$ explicit in the notation, suppose $\alpha_{2,t}(z_{t})=a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t})$ holds in a set of positive Lebesque measure in $z_{t}\in\mathbb{R}$. Since $\alpha_{2,t}(z_{t})$ and $\tilde{\alpha}_{2,t}(z_{t})$ are logistic, $\alpha_{2,t}(z_{t})$ and $a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t})$ are real analytic functions. Thus, by the Identity theorem, we have that $\alpha_{2,t}(z_{t})=a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t})$ for all $z_{t}$. Since this must hold for all $z_{t}$, it follows from $\underset{z_{t}\rightarrow-\infty}{\lim}\alpha_{2,t}(z_{t})=0$ and $\underset{z_{t}\rightarrow-\infty}{\lim}(a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t}))=a_{0,j}$ that $a_{0,j}=0$. Similarly, assuming $a_{1,j}\neq 0$, it follows from $a_{0,j}=0$, $\underset{z_{t}\rightarrow+\infty}{\lim}\alpha_{2,t}(z_{t})=1$, and $\underset{z_{t}\rightarrow+\infty}{\lim}(a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t}))=a_{0,j}+a_{1,j}$ that $a_{1,j}=1$. If $a_{1,j}=0$, then clearly $\alpha_{2,t}(z_{t})=a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t})$ holds only in the limit $z_{t}\rightarrow-\infty$, and thereby not in a set of positive Lebesgue measure. Thus, it must be that $a_{0,j}=0$ and $a_{1,j}=1$ in sets of positive Lebesgue measure, concluding that $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ almost everywhere in $z_{t}\in\mathbb{R}$.
[…]

### v5

@@ C.2 Proof of Lemma 2
To complete the first part of the proof, it needs to be shown that the diagonal elements of $D$ are either $1$ or 1-1. By the normalization of the unit variances of $e_{t}$ and $\tilde{e}_{t}$, $\text{Cov}(\tilde{e}_{t}|\mathcal{F}_{t-1})=\text{Cov}(Ce_{t}|\mathcal{F}_{t-1})=I_{d}$, which is equivalent to $C\text{Cov}(e_{t}|\mathcal{F}_{t-1})C^{\prime}=I_{d}$ and thereby $CC^{\prime}=I_{d}$. By substituting $C=DP$ to this expression, we obtain $DPP^{\prime}D^{\prime}=I_{d}$ and by the orthogonality of permutation matrix $DD^{\prime}=I_{d}$. That is, $d_{ii}^{2}=1$ for all i=1,,di=1,...,d, so $d_{ii}=\pm 1$. Since diagonal elements of $D$ are, hence, either $1$ or 1-1, it follows that $B_{y,t}$ is unique up to ordering of its columns and changing all signs in a column almost everywhere in $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
Logistic weights. Next, we show the ”moreover” part for the logistic transition weights defined in (2.3) with $M=2$ and endogenous $z_{t}$, in particular, a lagged endogenous variable obtained as $z_{t}=a^{\prime}\boldsymbol{y}_{t-1}$ with a fixed $a\neq 0$ (where $a$ has exactly one element equal to one and others equal to zero). Consider Equation (C.5) with two sets of parameters $\phi_{m},A_{m,1},...,A_{m,p}$, m=1,,Mm=1,...,M, $c,\gamma$ and $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, m=1,,Mm=1,...,M, $\tilde{c},\tilde{\gamma}$. The proof proceeds in the following steps: (1) we show that the AR parameters $\phi_{m},A_{m,1},...,A_{m,p}$, m=1,,Mm=1,...,M are identified (without relying on possible identification of the transition weights), and then, (2) given identification of the AR parameters, the identification of the transition weights and the parameters $c,\gamma$ is established.
Part 1. Denote $\boldsymbol{A}_{m}\equiv[A_{m,1}:...:A_{m,p}]$ $(d\times dp)$ and $\tilde{\boldsymbol{A}}_{m}\equiv[\tilde{A}_{m,1}:...:\tilde{A}_{m,p}]$ $(d\times dp)$, $m=1,2$. By making use of the identity $\alpha_{1,t}=1-\alpha_{2,t}$, write (C.5) as
$\begin{aligned} \alpha_{1,t}\big{[}(\boldsymbol{A}_{1}-\boldsymbol{A}_{2})\boldsymbol{y}_{t-1}+(\phi_{1}-\phi_{2})\big{]}&-\tilde{\alpha}_{1,t}\big{[}(\tilde{\boldsymbol{A}}_{1}-\tilde{\boldsymbol{A}}_{2})\boldsymbol{y}_{t-1}+(\tilde{\phi}_{1}-\tilde{\phi}_{2})\big{]}\\ &+\big{[}(\boldsymbol{A}_{2}-\tilde{\boldsymbol{A}}_{2})\boldsymbol{y}_{t-1}+(\phi_{2}-\tilde{\phi}_{2})\big{]}=0.\end{aligned}$ (C.7)
Since by assumption the transition weights depend only on the scalar switching variable $z_{t}=a^{\prime}\boldsymbol{y}_{t-1}$ (a single component of $\boldsymbol{y}_{t-1}$), we can analyze (C.7) on level sets of $z_{t}$ where the weights become constant. On each such set, Equation (C.7) reduces to an affine relation in $\boldsymbol{y}_{t-1}$ with fixed coefficients, which allows us to exploit linear-algebraic arguments to identify the AR matrices and intercepts. Because Equation (C.7) holds with probability one for $\boldsymbol{y}_{t-1}$, and Assumption 1 implies that the distribution of $\boldsymbol{y}_{t-1}$ has a density strictly positive almost everywhere in $\mathbb{R}^{dp}$, it follows that Equation (C.7) holds almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$. Below, we drop the time index from $\boldsymbol{y}_{t-1}$ and write simply $\boldsymbol{y}\in\mathbb{R}^{dp}$ for the generic lag vector when considering Equation (C.7) as a functional equation holding for almost every $\boldsymbol{y}\in\mathbb{R}^{dp}$.
For $v\in\mathbb{R}$, let $H_{v}\equiv\{\boldsymbol{y}\in\mathbb{R}^{dp}:a^{\prime}\boldsymbol{y}=v\}$ be the level set of $z_{t}$. Since (C.7) holds almost everywhere in $\boldsymbol{y}\in\mathbb{R}^{dp}$, it follows that for almost every value of $v=a^{\prime}\boldsymbol{y}$, (C.7) holds for almost every $\boldsymbol{y}\in H_{v}$. On $H_{v}$, the weights $\alpha_{1,t}=\alpha_{1}(v)$ and $\tilde{\alpha}_{1,t}=\tilde{\alpha}_{1}(v)$ are constants, so (C.7) becomes
$\begin{aligned} \alpha_{1}(v)\big{[}(\boldsymbol{A}_{1}-\boldsymbol{A}_{2})\boldsymbol{y}+(\phi_{1}-\phi_{2})\big{]}&-\tilde{\alpha}_{1}(v)\big{[}(\tilde{\boldsymbol{A}}_{1}-\tilde{\boldsymbol{A}}_{2})\boldsymbol{y}+(\tilde{\phi}_{1}-\tilde{\phi}_{2})\big{]}\\ &+(\boldsymbol{A}_{2}-\tilde{\boldsymbol{A}}_{2})\boldsymbol{y}+(\phi_{2}-\tilde{\phi}_{2})=0.\end{aligned}$ (C.8)
Since the left side of (C.8) is affine in $\boldsymbol{y}$, if it vanishes almost everywhere in $H_{v}$, it must vanish for all $\boldsymbol{y}\in H_{v}$. Thus, for almost every $v\in\mathbb{R}$, Equation (C.8) holds for all $\boldsymbol{y}\in H_{v}$, i.e., for all values of the dp1dp-1 free entries orthogonal to $a$.
To make the dependence on $v$ explicit, define
$L(v)$ $\equiv\alpha_{1}(v)(\boldsymbol{A}_{1}-\boldsymbol{A}_{2})-\tilde{\alpha}_{1}(v)(\tilde{\boldsymbol{A}}_{1}-\tilde{\boldsymbol{A}}_{2})+(\boldsymbol{A}_{2}-\tilde{\boldsymbol{A}}_{2}),$ (C.9)
$b(v)$ $\equiv\alpha_{1}(v)(\phi_{1}-\phi_{2})-\tilde{\alpha}_{1}(v)(\tilde{\phi}_{1}-\tilde{\phi}_{2})+(\phi_{2}-\tilde{\phi}_{2}),$ (C.10)
so that (C.8) reads $L(v)\boldsymbol{y}+b(v)=0$ for all $\boldsymbol{y}\in H_{v}$. Every $\boldsymbol{y}\in H_{v}$ can be written as
$\boldsymbol{y}=\boldsymbol{y}_{v}(v)+U\boldsymbol{y}_{\smallsetminus v},$ (C.11)
where $\boldsymbol{y}_{\smallsetminus v}\in\mathbb{R}^{dp-1}$, $U\in\mathbb{R}^{dp\times(dp-1)}$ has columns spanning $\mathcal{S}\equiv\{u:a^{\prime}u=0\}$ (so $a^{\prime}U=0$), and $\boldsymbol{y}_{v}(v)$ is any particular vector satisfying $a^{\prime}\boldsymbol{y}_{v}(v)=v$, e.g., $\boldsymbol{y}_{v}(v)\equiv\frac{v}{\|a\|^{2}}a$. Substituting gives, for almost every $v$,
$L(v)U\boldsymbol{y}_{\smallsetminus v}+(L(v)\boldsymbol{y}_{v}(v)+b(v))=0$ (C.12)
for all $\boldsymbol{y}_{\smallsetminus v}$. Hence, for almost every $v$, $L(v)U=0$ and $L(v)\boldsymbol{y}_{v}(v)+b(v)=0$. Each of entry of $L(v)U$ and $L(v)\boldsymbol{y}_{v}(v)+b(v)$ is a real-analytic function of $v$ (since $\alpha_{1}(v),\tilde{\alpha}_{1}(v)$ are logistic and hence real-analytic, and $\boldsymbol{y}_{v}(v)$ is affine in $v$). It follows that by the Identity Theorem, both $L(v)U=0$ and $L(v)\boldsymbol{y}_{v}(v)+b(v)=0$ hold for all $v\in\mathbb{R}$. Consequently, (C.8) holds for all $v$ and $\boldsymbol{y}\in H_{v}$.
Subtracting (C.8) for two (distinct) points $\boldsymbol{y}^{(1)},\boldsymbol{y}^{(2)}\in H_{v}$ eliminates the intercepts, giving
$L(v)u=0\quad\text{for all }u\equiv\boldsymbol{y}^{(1)}-\boldsymbol{y}^{(2)}\in\mathcal{S},$ (C.13)
where $\mathcal{S}=\{u:a^{\prime}u=0\}$ is the (dp1)(dp-1)-dimensional subspace of vectors orthogonal to $a$, i.e., the set of displacements in $\mathbb{R}^{dp}$ that keep $a^{\prime}\boldsymbol{y}$ (and hence $z_{t}$) fixed. Note that the difference $u$ does not depend on the level $v$ of switching variable $z_{t}$, since the related component of $u$ is always zero by construction.
Now, observe that if $M\in\mathbb{R}^{d\times dp}$ satisfies $Mu=0$ for all $u\in\mathcal{S}$, then there exists a unique $\beta\in\mathbb{R}^{d}$ such that [^11] To see this, pick any $w$ with $a^{\prime}w=1$ and decompose $\boldsymbol{y}=u+(a^{\prime}\boldsymbol{y})w$ with $u\in\mathcal{S}$. Then $M\boldsymbol{y}=(a^{\prime}\boldsymbol{y})Mw$, and setting $\beta\equiv Mw$ gives $M\boldsymbol{y}=\beta a^{\prime}\boldsymbol{y}$ for all $\boldsymbol{y}$, i.e., $M=\beta a^{\prime}$. To show uniqueness, suppose $M=\tilde{\beta}a^{\prime}$ holds also for some $\tilde{\beta}\in\mathbb{R}^{d}$. Then, $(\beta-\tilde{\beta})a^{\prime}=0$, and multiplying this by $w$ from the right gives $\beta-\tilde{\beta}=Mw-\tilde{\beta}=0$, showing $\tilde{\beta}=Mw$ and hence $\tilde{\beta}=\beta$.
$M\boldsymbol{y}=\beta a^{\prime}\boldsymbol{y}\ \ \text{for all}\ \ \boldsymbol{y}.$ (C.14)
For any $u\in\mathcal{S}$, Equation (C.13) can be expressed as
$\alpha_{1}(v)(\boldsymbol{A}_{1}-\boldsymbol{A}_{2})u-\tilde{\alpha}_{1}(v)(\tilde{\boldsymbol{A}}_{1}-\tilde{\boldsymbol{A}}_{2})u+(\boldsymbol{A}_{2}-\tilde{\boldsymbol{A}}_{2})u=0.$ (C.15)
For each component k=1,,dk=1,...,d, define
$u^{(1)}_{k}\equiv[(\boldsymbol{A}_{1}-\boldsymbol{A}_{2})u]_{k},\quad u^{(2)}_{k}\equiv[(\tilde{\boldsymbol{A}}_{1}-\tilde{\boldsymbol{A}}_{2})u]_{k},\quad u^{(0)}_{k}\equiv[(\boldsymbol{A}_{2}-\tilde{\boldsymbol{A}}_{2})u]_{k}.$ (C.16)
Then, the $k$th component of (C.15) is
$\alpha_{1}(v)u^{(1)}_{k}-\tilde{\alpha}_{1}(v)u^{(2)}_{k}+u^{(0)}_{k}=0\quad\forall v\in\mathbb{R}.$ (C.17)
Let $r(v)\equiv(\alpha_{1}(v),-\tilde{\alpha}_{1}(v),1)\in\mathbb{R}^{3}$ and $w_{k}\equiv(u^{(1)}_{k},u^{(2)}_{k},u^{(0)}_{k})\in\mathbb{R}^{3}$, so that (C.17) can be expressed as $r(v)^{\prime}w_{k}=0$, i.e., $w_{k}$ is orthogonal to the set $\{r(v):v\in\mathbb{R}\}$.
If $\alpha_{1}(v)=\tilde{\alpha}_{1}(v)$ for all $v$, then $r(v)=(\alpha_{1}(v),-\alpha_{1}(v),1)$, so the span of $r(v)$ is only $2$-dimensional. Then, evaluating (C.17) at two distinct values of $v$ yields linear conditions that force $w_{k}=0$. If $\alpha_{1}(v)=\tilde{\alpha}_{1}(v)$ does not hold identically, the set $\{r(v):v\in\mathbb{R}\}$ spans the whole $\mathbb{R}^{3}$. [^12] If not, then there exists a vector $(c_{1},c_{2},c_{3})\neq 0$ with $c_{1}\alpha_{1}(v)-c_{2}\tilde{\alpha}_{1}(v)+c_{3}=0$ for all $v$. Taking the limit $v\rightarrow\infty$ yields $c_{3}=0$, as $\underset{v\rightarrow\infty}{\lim}\alpha_{1}(v)=\underset{v\rightarrow\infty}{\lim}\tilde{\alpha}_{1}(v)=0$. Then $c_{1}\alpha_{1}(v)=c_{2}\tilde{\alpha}_{1}(v)$ for all $v$, and taking the limit $v\rightarrow$-$\infty$ yields $c_{1}=c_{2}$, as $\underset{v\rightarrow-\infty}{\lim}\alpha_{1}(v)=\underset{v\rightarrow-\infty}{\lim}\tilde{\alpha}_{1}(v)=1$. This forces $\alpha_{1}(v)=\tilde{\alpha}_{1}(v)$, which is a contradiction. Therefore the only vector $w_{k}$ orthogonal to all $r(v)$ is the zero vector, so $w_{k}=0$.
Since $w_{k}=0$ for every k=1,,dk=1,...,d, we have
$(\boldsymbol{A}_{1}-\boldsymbol{A}_{2})u=0,\ \ (\tilde{\boldsymbol{A}}_{1}-\tilde{\boldsymbol{A}}_{2})u=0,\ \ \text{and}\ \ (\boldsymbol{A}_{2}-\tilde{\boldsymbol{A}}_{2})u=0$ (C.18)
for all $u\in\mathcal{S}$. Applying the result given in (C.14) then shows
$(\boldsymbol{A}_{1}-\boldsymbol{A}_{2})\boldsymbol{y}=\beta_{1}a^{\prime}\boldsymbol{y},\ \ (\tilde{\boldsymbol{A}}_{1}-\tilde{\boldsymbol{A}}_{2})\boldsymbol{y}=\beta_{2}a^{\prime}\boldsymbol{y},\ \ \text{and}\ \ (\boldsymbol{A}_{2}-\tilde{\boldsymbol{A}}_{2})\boldsymbol{y}=\beta_{0}a^{\prime}\boldsymbol{y}$ (C.19)
for some $\beta_{0},\beta_{1},\beta_{2}\in\mathbb{R}^{d}$ and all $\boldsymbol{y}$. Substituting these identities into (C.8) and taking $\boldsymbol{y}\in H_{v}$ (so $a^{\prime}\boldsymbol{y}=v$) gives
$\alpha_{1}(v)(\beta_{1}v+\phi_{1}-\phi_{2})-\tilde{\alpha}_{1}(v)(\beta_{2}v+\tilde{\phi}_{1}-\tilde{\phi}_{2})+(\beta_{0}v+\phi_{2}-\tilde{\phi}_{2})=0$ (C.20)
for all $v$.
For $v\neq 0$, dividing (C.20) by $v$ and taking the limit $v\rightarrow\infty$ yields $\beta_{0}=0$, as $\underset{v\rightarrow\infty}{\lim}\alpha_{1}(v)=\underset{v\rightarrow\infty}{\lim}\tilde{\alpha}_{1}(v)=0$. Plugging this into the last equation in (C.19) gives $(\boldsymbol{A}_{2}-\tilde{\boldsymbol{A}}_{2})\boldsymbol{y}=0$ for all $\boldsymbol{y}$, and thus, $\boldsymbol{A}_{2}=\tilde{\boldsymbol{A}}_{2}$. With $\beta_{0}=0$, dividing (C.20) by $v$ and taking the limit $v\rightarrow$-$\infty$ yields $\beta_{1}=\beta_{2}$, as $\underset{v\rightarrow-\infty}{\lim}\alpha_{1}(v)=\underset{v\rightarrow-\infty}{\lim}\tilde{\alpha}_{1}(v)=1$. Plugging this and $\boldsymbol{A}_{2}=\tilde{\boldsymbol{A}}_{2}$ in the first two equations in (C.19) yields $(\boldsymbol{A}_{1}-\tilde{\boldsymbol{A}}_{1})\boldsymbol{y}=0$ for all $\boldsymbol{y}$, and hence, $\boldsymbol{A}_{1}=\tilde{\boldsymbol{A}}_{1}$.
Also, $\alpha_{1}(v)v\rightarrow 0$ as $v\rightarrow\infty$, since for $v>0$ we have $0\leq v\alpha_{1}(v)\leq ve^{-\gamma(v-c)}=\frac{v}{e^{\gamma v}}e^{\gamma c}\rightarrow 0$ as $v\rightarrow\infty$ by L’Hôpital’s rule. Consequently, with $\beta_{0}=0$, taking the limit $v\rightarrow\infty$ in (C.20) gives $\phi_{2}=\tilde{\phi}_{2}$. Since the denominator of $v(\alpha_{1}(v)-\tilde{\alpha}_{1}(v))=\frac{v\exp\{\tilde{\gamma}(v-\tilde{c})\}-v\exp\{\gamma(v-c)\}}{(1+\exp\{\gamma(v-c)\})(1+\exp\{\tilde{\gamma}(v-\tilde{c})\})}$ tends to one as $v\rightarrow$-$\infty$, it follows that $\underset{v\rightarrow-\infty}{\lim}v(\alpha_{1}(v)-\tilde{\alpha}_{1}(v))=e^{-\tilde{\gamma}\tilde{c}}\underset{v\rightarrow-\infty}{\lim}ve^{\tilde{\gamma}v}-e^{-\gamma c}\underset{v\rightarrow-\infty}{\lim}ve^{\gamma v}=0$ by applying L’Hôpital’s rule. Therefore, with $\beta_{0}=0$, $\beta_{1}=\beta_{2}$, and $\phi_{2}=\tilde{\phi}_{2}$, taking the limit $v\rightarrow$-$\infty$ in (C.20) gives $\phi_{1}=\tilde{\phi}_{1}$. Thus, the AR parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,2$, are uniquely identified.
Part 2. Given identification of $\phi_{m}$ and $A_{m,i}$, we can substitute $\phi_{m}=\tilde{\phi}_{m}$ and $A_{m,i}=\tilde{A}_{m,i}$, $m=1,2$, i=1,,pi=1,...,p, to Equation (C.8) to obtain that, for every $v\in\mathbb{R}$,
$(\alpha_{1}(v)-\tilde{\alpha}_{1}(v))\big{[}\mu_{1}(\boldsymbol{y})-\mu_{2}(\boldsymbol{y})\big{]}=0\ \ \text{for all}\ \ \boldsymbol{y}\in H_{v}=\{\boldsymbol{y}\in\mathbb{R}^{dp}:a^{\prime}\boldsymbol{y}=v\},$ (C.21)
where $\mu_{m}(\boldsymbol{y})\equiv\phi_{m}+\boldsymbol{A}_{m}\boldsymbol{y}$, $m=1,2$. Hence, for each fixed $v$, either $\alpha_{1}(v)=\tilde{\alpha}_{1}(v)$ or $\mu_{1}(\boldsymbol{y})=\mu_{2}(\boldsymbol{y})$ for all $\boldsymbol{y}\in H_{v}$. Considering the latter case, the equality $\mu_{1}(\boldsymbol{y})=\mu_{2}(\boldsymbol{y})$ can be written as
$(\phi_{1}-\phi_{2})+(\boldsymbol{A}_{1}-\boldsymbol{A}_{2})\boldsymbol{y}=0.$ (C.22)
Subtracting for two (distinct) points $\boldsymbol{y}^{(1)},\boldsymbol{y}^{(2)}\in H_{v}$ eliminates the intercepts and gives
$(\boldsymbol{A}_{1}-\boldsymbol{A}_{2})u=0\ \ \text{for all}\ \ \boldsymbol{u}=\boldsymbol{y}^{(1)}-\boldsymbol{y}^{(2)}\in\mathcal{S},$ (C.23)
where $\mathcal{S}=\{u:a^{\prime}u=0\}$ as before. Then, applying (C.14) with $M=\boldsymbol{A}_{1}-\boldsymbol{A}_{2}$ yields
$(\boldsymbol{A}_{1}-\boldsymbol{A}_{2})\boldsymbol{y}=\beta_{5}a^{\prime}\boldsymbol{y}$ (C.24)
for some $\beta_{5}\in\mathbb{R}^{d}$ and all $\boldsymbol{y}$. Substituting this to (C.22) and taking $\boldsymbol{y}\in H_{v}$ so that $a^{\prime}\boldsymbol{y}=v$ implies $(\phi_{1}-\phi_{2})+\beta_{5}v=0.$
Thus, if $\beta_{5}\neq 0$, the equality $\mu_{1}(\boldsymbol{y})=\mu_{2}(\boldsymbol{y})$ can hold only at a single level $v_{0}$ that satisfies $(\phi_{1}-\phi_{2})+\beta_{5}v_{0}=0$. If $\beta_{5}=0$, it follows from (C.24) that $\boldsymbol{A}_{1}=\boldsymbol{A}_{2}$ and thus from (C.22) that also $\phi_{1}=\phi_{2}$, contradicting Condition (1) of Lemma 2. Therefore, it must be that $\alpha_{1}(v)=\tilde{\alpha}_{1}(v)$ for all $v\neq v_{0}$. However, since $\alpha_{1}(v),\tilde{\alpha}_{1}(v)$ are real-analytic in $v$ and coincide on $\mathbb{R}\setminus\{v_{0}\}$ that has a positive Lebesgue measure, they must agree everywhere by the Identity Theorem, concluding that $\alpha_{1,t}=\tilde{\alpha}_{1,t}$ identically.
[…]